# OSINT Clinic: Co-designing AI-Augmented Collaborative OSINT Investigations for Vulnerability Assessment

### v2

@@ 3.1. Study Protocol
Figure 1. Study Diagram: (1) Study 1 was formative and aimed at identifying the key challenges in the Intelligence Cycle phases of OSINT investigations. We asked participants to reflect on previous investigations and document the steps and challenges faced during Design Workshop 1. (2) In Study 2, we introduced generative AI probes in Design Workshops 2–4 to explore how AI could address challenges experienced earlier. During the workshops, we identified areas where generative AI was effective and also its limitations. We focused on the emergent challenges of collaboration and leadership during design workshops 4 and 5, with a collaborative AI platform (Team-GPT) as a probe. (3) In Study 3, we piloted OSINT-based vulnerability assessments with three small businesses, integrating curated generative AI prompts into ongoing investigations. This real-world application highlighted the practical utility and surfaced challenges of using foundational generative AI models like ChatGPT in vulnerability assessments, especially in collaborative settings.
@@
Our longitudinal study spanned two semesters. Study 1 (Design Workshop 1) was conducted after three months of OSINT training. Right after, we ran Study 2 (Design workshops 2–6) over two months during the end of the first semester and the beginning of the second. Finally, Study 3 involved two focus group interviews conducted over a span of one month at the end of the second semester. Figure 1 represents the flow of the research study with the objective, data source(s), and probe(s) for each of Studies 1-3.
@@ 3.2. Participant Recruitment
$3.$ Limitations
Our study involved only six participants over two semesters, which may limit the diversity of needs and design ideas explored. Previous research in intelligence analysis, including studies with 2–10 participants (Kang and Stasko, 2014; Chin Jr et al., 2009; Riebe et al., 2023), often face similar constraints due to the sensitive nature of the work and client involvement.
In Study 2, short sessions (10–20 minutes) for trying out prompts and features may have restricted idea generation. Future studies could expand on ideas by including longer workshops with power users of generative AI tools or OSINT practitioners.
Generative AI probes can produce incorrect responses or ”hallucinations,” potentially limiting the reliability of disseminated results. While our focus on verification and cross-referencing addresses some issues, further research is needed to develop accountable human-AI collaboration frameworks (Chen et al., 2023).
4. Study 1: Identifying Challenges with OSINT Tasks
@@ 4.2. Analysis
The first author transcribed the recorded workshop and, in collaboration with the rest of the research team, conducted an inductive thematic analysis (Braun and Clarke, 2006) of the transcript to identify the challenges. We aligned the identified challenges with the five phases of the OSINT Intelligence Model (OIM) presented by Hwang et al. (Hwang et al., 2022) (see Figure 2), which closely follows the IC. We denote the first stage of “identifying sources for purpose” as “planning” to match the IC, as it entails the same task of figuring “where and how to get information.”
@@ 4.3.5. Dissemination
5. Study 2: Using AI Probes to Explore Design Ideas for OSINT Tasks
@@ 5.2.1. Planning
P3 and P5 suggested that ChatGPT should provide a rating or value scale to help users better understand the reputation of a tool or source before using it. Participants wanted AI to offer unbiased views on privacy and trust, with P6 stating, “I would like AI to be able to provide comments on the privacy and amount of trust I can place in different verification and informative sources.” Participants explored web browsing and custom GPT models and thought they could enhance operational security training and simulations. They also noted the need for regular updates and specificity in the AI’s knowledge base due to the rapid pace of change in the OSINT tools ecosystem.
@@
Participants found generative AI to be effective in assisting with task division and resource identification, making it easier to manage team-based OSINT investigations. P5 reflected, ”I basically just asked it to come up with a list of subtasks and a list of tools and techniques…and it worked perfectly”. However, there were challenges with the AI’s limitations in providing specific details related to smaller businesses as P4 observed, “ChatGPT provides little to no information on smaller businesses”. There were also difficulties crafting prompts to avoid “guardrails” against inappropriate content or eliciting new, non-repetitive information.
One participant idea to improve the experience was for ChatGPT to link directly to tools and resources it recommends, simplifying the follow-up steps in investigations. They saw value in ChatGPT’s ability to create visually appealing documents and diagrams, which would aid in team discussions and presenting findings. The use of advanced features like custom GPT models trained with OSINT-specific databases and plugins that create visual templates were also proposed as ways to make the AI more tailored to the unique needs of OSINT investigations.
@@ 5.2.2. Collection
Participants found ChatGPT to be particularly effective for discovering resources when they asked for specific tools. They could generate alternative options through follow-ups when the initial suggestions did not meet their needs. However, challenges were noted, particularly regarding the AI’s outdated knowledge base, as it only contained information on tools up to January 2022. Furthermore, the tools suggested were not always available. P1 described the process of verifying each tool suggested by ChatGPT: “You have to painstakingly copy and paste each one [tool] and research it yourself.” The need to manually verify AI’s suggestions added to the complexity of using these tools effectively.
A common suggestion for improvement was for ChatGPT to provide more contextual information on an OSINT tool, or as P6 explained: “getting a comprehensible description/summary of the tool so I know what and how to use it in my work.” Participants wanted more clarity about the tool’s data sources, whether tools had a paywall, as well as how to find specific details about their trustworthiness. However, they also noted that while ChatGPT’s web browsing capabilities could cover a wide range of knowledge, the credibility of the sources provided was questionable and emphasized the need for careful cross-verification.
@@
Participants particularly appreciated the AI’s ability to quickly lay out different sections and specify what should go under each. Participants faced challenges in reformatting the AI-generated templates to make them more presentable. P1 explained, “ChatGPT is great for getting your data quickly into format but honestly it doesn’t look very good. With a little bit of editing and using a format made by a human I think it would look much better.” Importantly, the AI’s responses varied depending on the prompts, leading to inconsistencies in the templates generated across different users that impeded collaboration.
Participants wanted ChatGPT to generate files containing templates, such as spreadsheets or documents, which would reduce the need for additional formatting. P6 recalled, “I tried asking ChatGPT to make the template into a spreadsheet, but it is unable to directly make files. It would be more helpful if ChatGPT could generate files.” P3 highlighted the potential of using plugins to generate and convert templates more efficiently. Other participants proposed the idea idea of a prompt library, where successful prompts for templates could be shared and rated by others.
@@ 5.2.3. Processing
Participants reiterated the need to directly generate files in the specified format, such as spreadsheets with particular columns, to reduce manual formatting efforts. They explored the integration of automation tools like web scraping through Python libraries, though they found it frustrating that ChatGPT could not perform these tasks directly and instead provided instructions for them. These suggestions highlight the potential for ChatGPT to be more automated in the data organization and documentation process.
@@
Participants noted that ChatGPT was useful in identifying potential verification sources and organizing data. However, tool suggestions often fell short of providing the detailed, specific information needed to trust the AI’s recommendations, as sources or tools were unfamiliar.
Participants suggested a more iterative experience for diving deeper. P1 said, ”I would like AI to be more interactive and help complete the investigation with me by giving recommendations on places to find information not found”. They wanted detailed ratings or community sentiment analysis of tools for reliability. P2 also cautioned about relying too heavily on AI’s web browsing capabilities, as they still needed to be double-checked for accuracy.
@@ 5.2.4. Analysis
Participants found ChatGPT’s explanations to be somewhat helpful. As P3 noted, “It was useful for elaborating on vulnerabilities and explaining in simpler terms…especially for a non-technical audience.”. However, P6 found output adjustments to be hard, ”Using ChatGPT for interpretation requires specific prompts for better detail. It can be helpful, but you need to be clear about what you’re asking for”. P5 found that ChatGPT did not provide detailed security analysis when using web browsing capabilities, often citing limitations in its scope. P3 mentioned specialized plugins like Check Point Security, which evaluates the security status of IP addresses, URLs, and file hashes, to augment ChatGPT’s general responses.
@@
ChatGPT helped present supporting data for reports and generate recommendations. Participants found it difficult to identify the most relevant actions from a large number of recommendations. Similar to previous challenges, ChatGPT and custom GPTs often failed to generate specific and actionable recommendations tailored to the unique needs of a given situation, sometimes refusing a detailed analysis for being out of scope or unethical. Participants wanted support in creating data visualizations for recommendations, such as Gantt charts, to better illustrate timelines and implementation strategies.
@@ 5.2.5. Dissemination
Participants found that using ChatGPT to simulate client feedback was particularly useful in preparing for client interactions. P1 emphasized that iterative feedback from ChatGPT helped in making the recommendations more actionable and better tailored to client needs. However, P4 observed that the iterative process was often limited to removing, or reformatting content rather than delving into more detailed analysis. Participants also worried about privacy issues in involving the AI at this final stage, when security flaws were at their most distilled and contextualized. Some suggested that turning off chat history or using personal chat configurations could enhance privacy during the iterative process of sharing sensitive information like vulnerabilities.
@@ 5.3.2. Results
In conclusion, Study 2 explored how generative AI could address key challenges in OSINT investigations and developed design ideas. Building on the insights, participants used generative AI in conducting OSINT-based vulnerability assessments on small businesses during Study 3.
@@ 6.1. Method
At an initial meeting, the researchers described the process to the clients, set expectations around timelines and their involvement, and answered questions. Clients were asked to select from a menu of services based on the categories mentioned in Table 1. After the meeting, if clients wanted to proceed, they signed a letter of agreement consenting to being investigated and provided some basic information about their digital infrastructure.
As mentioned in Studies 1 and 2, participants had training and practice with OSINT tradecraft for over a semester, as well as experience using generative AI platforms like ChatGPT and Team-GPT. Participants worked together as a single team of six members for the first investigation, and then as teams of 2–3 members each (including a designated leader) for the next two, which ran in parallel. The teams spent around 2.5 weeks per investigation, from data collection and analysis to developing a presentation describing the vulnerabilities identified and making recommendations to mitigate them.
@@
We followed best practices for security, privacy, and ethics established in previous research on student-led real-world OSINT investigations (Mukhopadhyay et al., 2024; cli, 2024). The students signed Non-Disclosure Agreements (NDAs) to protect the confidentiality of their findings except for sharing with the customer. The students also enhanced their safety by taking operational security measures like using Virtual Private Networks (VPNs), Virtual Machines (VMs), role-based access control and encryption for storing and sharing data, and sock puppet accounts.
@@
6.2. Analysis
First, we performed deductive thematic analysis (Braun and Clarke, 2006) on the focus group transcripts to understand (1) the use of generative AI, and (2) the challenges that emerged and how participants circumvented them. We also analyzed the shared chats on the Team-GPT workspace to monitor the use of the prompt library. Then, we aggregated the three client evaluations of vulnerability assessments to understand their utility.
@@ 6.3. Results
Utility of generative AI during investigations
@@
Challenges and limitations faced during investigations
@@
Generative AI in the form of a foundational model like ChatGPT, showed limitations in data collection and analysis. A recurring theme in Study 2, participants again found that AI often provided broad, generalized responses that required further refinement, as P5 mentioned, ”It tends to give broad responses and doesn’t always provide the specific details needed for verification”. This may be partly due to the use of abstracted prompts to bypass privacy issues. Consequently, participants spent the majority of their investigation outside the Team-GPT or ChatGPT platforms, using external tools and searching websites or social media accounts.
@@
Usefulness of vulnerability assessments based on client survey
@@ 7.1.1. Adapting the method for OSINT investigations
We adapted and extended the Matchmaking for AI method (Liu et al., 2024) for an OSINT context to structure design workshops and generate innovative yet feasible ideas. We followed the three key steps of the original method. First, for mapping stakeholder domain expertise and AI needs, we focused on challenges and learning goals across the OSINT Intelligence Model for OSINT investigations. For Step 2, playing with AI probes to make participants familiar with “affordances of the technology,” we used ChatGPT and Team-GPT as probes. For the third step, brainstorming AI design elements, we identified the strengths and limitations of general-purpose LLMs and explored ways to enhance the user experience.
However, we also extended this method by addressing two limitations revealed by our study context. First, our brainstorming sessions in a group setting, with a mix of co-located and online setups alleviated the limits of interpersonal engagement and creative expression among participants in a remote, individual study setting. Second, we used ChatGPT and Team-GPT as functional technology probes (Hutchinson et al., 2003) from the start, compared to Wizard-of-Oz simulations, to help participants envision more diverse and concrete AI solutions.
@@ 7.1.2. Extending the method for collaboration and learning
Chin Jr. et al. highlight collaborative analysis as a critical approach for intelligence analysis, and we report on its evolution over time by observing AI interactions and teamwork across workshops and real investigations (Chin Jr et al., 2009). In Study 2 (Section 5.3), we investigate new challenges and design ideas that emerge as the team collaborates and incorporates generative AI tools over time. We propose a non-linear design process, compared to the linear structure in the original method, to address collaboration and leadership challenges. This cyclical structure promotes fluidity between the phases of co-design through subsequent workshops based on results from previous ones.

### v3

@@ 3.1. Study Protocol
Figure 1. Study Diagram: (1) Study 1 was formative and aimed at identifying the key challenges in the Intelligence Cycle phases of OSINT investigations. We asked participants to reflect on previous investigations and document the steps and challenges faced during Design Workshop 1. (2) In Study 2, we introduced generative AI probes in Design Workshops 2–4 to explore how AI could address challenges experienced earlier. During the workshops, we identified areas where generative AI was effective and also its limitations. We focused on the emergent challenges of collaboration and leadership during Design Workshops 5 and 6, with a collaborative AI platform (Team-GPT) as a probe. (3) In Study 3, we piloted OSINT-based vulnerability assessments with three small businesses, integrating curated generative AI prompts into ongoing investigations. This real-world application highlighted the practical utility and surfaced challenges of using foundational generative AI models like ChatGPT in vulnerability assessments, especially in collaborative settings.
@@
Our longitudinal study spanned two semesters. Study 1 (Design Workshop 1) was conducted after three months of OSINT training. Right after, we ran Study 2 (Design workshops 2–6) over two months during the end of the first semester and the beginning of the second. Finally, Study 3 involved two focus group interviews conducted over a span of one month at the end of the second semester. The longitudinal design of this study was essential for us to observe how the use of generative AI evolved over time and across different phases of the Intelligence Cycle. This extended engagement provided insights into how AI is integrated, how collaboration challenges develop, and how leadership roles evolve in response to AI support. Figure 1 represents the flow of the research study with the objective, data source(s), and probe(s) for each of Studies 1-3.
@@ 3.2. Participant Recruitment
4. Study 1: Identifying Challenges and Learning Goals with OSINT Tasks
@@ 4.2. Analysis
The first author transcribed the recorded workshop and, in collaboration with the rest of the research team, conducted an inductive thematic analysis (Braun and Clarke, 2006) of the transcript to identify the challenges. We aligned the identified challenges with the five phases of the OSINT Intelligence Model (OIM) presented by Hwang et al. (Hwang et al., 2022) (see Figure 2), which closely follows the IC. Hwang et al’s model focuses on continuous iteration and practical applications to cybersecurity (Hwang et al., 2022). We denote the first stage of “identifying sources for purpose” as “planning” to match the IC, as it entails the same task of figuring “where and how to get information.”
@@ 4.3.5. Dissemination
5. Study 2: Using Generative AI Probes to Achieve Learning Goals for OSINT Investigation
@@ 5.2.1. Planning
Participants found generative AI to be effective in assisting with task division and resource identification, making it easier to manage team-based OSINT investigations. P5 reflected, ”I basically just asked it to come up with a list of subtasks and a list of tools and techniques…and it worked perfectly”. However, there were challenges with the AI’s limitations in providing specific details related to smaller businesses as P4 observed, “ChatGPT provides little to no information on smaller businesses”. There were also difficulties crafting prompts to avoid “guardrails” against inappropriate content or eliciting new, non-repetitive information. Participants proposed custom GPT models trained with OSINT-specific databases and plugins to make the AI more tailored to the unique needs of OSINT investigations.
@@ 5.2.2. Collection
Participants found ChatGPT to be particularly effective for discovering resources when they asked for specific tools. They could generate alternative options through follow-ups when the initial suggestions did not meet their needs. However, challenges were noted, particularly regarding the AI’s outdated knowledge base, as it only contained information on tools up to January 2022. Furthermore, the tools suggested were not always available. P1 described the process of verifying each tool suggested by ChatGPT: “You have to painstakingly copy and paste each one [tool] and research it yourself.” Participants noted that while ChatGPT’s web browsing capabilities could cover a wide range of knowledge, the credibility of the sources provided was questionable and emphasized the need for careful cross-verification.
@@
Participants particularly appreciated the AI’s ability to quickly lay out different sections and specify what should go under each. Participants faced challenges in reformatting the AI-generated templates to make them more presentable. P1 explained, “ChatGPT is great for getting your data quickly into format but honestly it doesn’t look very good. With a little bit of editing and using a format made by a human I think it would look much better.” Importantly, the AI’s responses varied depending on the prompts, leading to inconsistencies in the templates generated across different users that impeded collaboration.Participants proposed the idea idea of a prompt library, where successful prompts for templates could be shared and rated by others.
@@ 5.2.3. Processing
Participants noted that ChatGPT was useful in identifying potential verification sources and organizing data. However, tool suggestions often fell short of providing the detailed, specific information needed to trust the AI’s recommendations, as sources or tools were unfamiliar. P2 cautioned about relying too heavily on AI’s web browsing capabilities, as they still needed to be double-checked for accuracy.
@@ 5.2.4. Analysis
Participants found ChatGPT’s explanations to be somewhat helpful. As P3 noted, “It was useful for elaborating on vulnerabilities and explaining in simpler terms…especially for a non-technical audience.”. However, P6 found output adjustments to be hard, ”Using ChatGPT for interpretation requires specific prompts for better detail. It can be helpful, but you need to be clear about what you’re asking for”. P5 found that ChatGPT did not provide detailed security analysis when using web browsing capabilities, often citing limitations in its scope.
@@
ChatGPT helped present supporting data for reports and generate recommendations. Participants found it difficult to identify the most relevant actions from a large number of recommendations. Similar to previous challenges, ChatGPT and custom GPTs often failed to generate specific and actionable recommendations tailored to the unique needs of a given situation, sometimes refusing a detailed analysis for being out of scope or unethical.
@@ 5.2.5. Dissemination
Participants found that using ChatGPT to simulate client feedback was particularly useful in preparing for client interactions. P1 emphasized that iterative feedback from ChatGPT helped in making the recommendations more actionable and better tailored to client needs. However, P4 observed that the iterative process was often limited to removing, or reformatting content rather than delving into more detailed analysis. Participants also worried about privacy issues in involving the AI at this final stage, when security flaws were at their most distilled and contextualized.
@@ 5.3.2. Results
In conclusion, Study 2 explored how generative AI could address key challenges in OSINT investigations and developed design ideas. The findings show how AI supported learning and helped achieve the learning goals. Building on the insights from using it in different parts of the investigation, participants used generative AI in conducting OSINT-based vulnerability assessments on small businesses during Study 3.
@@ 6.1. Method
We structured the consultation process using the Understand-Investigate-Advise (UIA) framework introduced by Havron et al.(Havron et al., 2019; Freed et al., 2019). In the initial Understand phase, the researchers met with the clients to explain the consultation process, set expectations regarding timelines and involvement, and answer any questions. Clients were then asked to select relevant services from a menu of tasks based on the categories outlined in Table1. Those who wished to proceed signed a letter of agreement consenting to the investigation and provided basic information about their digital infrastructure.
In the Investigate phase, participants drew on their semester-long training in OSINT tradecraft and their experience with generative AI platforms like ChatGPT and Team-GPT (as discussed in Studies 1 and 2). For the first investigation, all six participants collaborated as a single team. In subsequent investigations, they split into smaller teams of 2–3 members, each with a designated leader. Over approximately 2.5 weeks per investigation, the teams collected and analyzed data, identifying potential vulnerabilities in the clients’ digital systems.
Finally, during the Advise phase, the teams synthesized their findings into presentations. These presentations detailed the identified vulnerabilities and offered tailored recommendations for mitigation. By delivering clear and actionable advice, the teams ensured clients received practical strategies to improve their cybersecurity posture.
@@
We followed best practices for security, privacy, and ethics established in previous research on student-led real-world OSINT investigations (Mukhopadhyay et al., 2024; cli, 2024a). The students signed Non-Disclosure Agreements (NDAs) to protect the confidentiality of their findings except for sharing with the customer. The students also enhanced their safety by taking operational security measures like using Virtual Private Networks (VPNs), Virtual Machines (VMs), role-based access control and encryption for storing and sharing data, and sock puppet accounts.
@@
In terms of analysis, first, we performed deductive thematic analysis (Braun and Clarke, 2006) on the focus group transcripts to understand (1) the use of generative AI, and (2) the challenges that emerged and how participants circumvented them. We also analyzed the shared chats on the Team-GPT workspace to monitor the use of the prompt library. Then, we aggregated the three client evaluations of vulnerability assessments to understand their utility.
@@ 6.2. Results
6.2.1. Utility of generative AI during investigations
@@
6.2.2. Challenges and limitations faced during investigations
@@
Generative AI in the form of a large language model like ChatGPT, showed limitations in data collection and analysis. A recurring theme in Study 2, participants again found that AI often provided broad, generalized responses that required further refinement, as P5 mentioned, ”It tends to give broad responses and doesn’t always provide the specific details needed for verification”. This may be partly due to the use of abstracted prompts to bypass privacy issues. Consequently, participants spent the majority of their investigation outside the Team-GPT or ChatGPT platforms, using external tools and searching websites or social media accounts.
@@
6.2.3. Usefulness of vulnerability assessments based on client survey
@@ 7.1.1. Adapting the method for OSINT investigations
We adapted and extended the Matchmaking for AI method (Liu et al., 2024) for an OSINT context to structure design workshops and generate innovative yet feasible ideas. We followed the three key steps of the original method. First, for mapping stakeholder domain expertise and AI needs, we focused on challenges and learning goals across the OSINT Intelligence Model for OSINT investigations. For Step 2, playing with AI probes to make participants familiar with “affordances of the technology,” we used ChatGPT and Team-GPT as probes. For the third step, brainstorming AI design elements, we identified ways to tailor general-purpose LLMs in developing and applying OSINT skills.
We also extended this method by addressing two limitations revealed by our study context. First, our brainstorming sessions in a group setting, with a mix of co-located and online setups alleviated the limits of interpersonal engagement and creative expression among participants in a remote, individual study setting. Second, we used ChatGPT and Team-GPT as functional technology probes (Hutchinson et al., 2003), compared to Wizard-of-Oz simulations, to help participants envision more diverse and concrete AI solutions.
@@ 7.1.2. Extending the method for collaboration and learning
Chin Jr. et al. highlight collaborative analysis as a critical approach for intelligence analysis, and we report on its evolution over time by observing AI interactions and teamwork across workshops and real investigations (Chin Jr et al., 2009). In Study 2 (Section 5.3), we investigate new challenges and design ideas that emerge as the team collaborates and incorporates generative AI tools over time. Our longitudinal and cyclical study design promotes fluidity between the phases of co-design and helps to study collaboration.